OpenStudy (anonymous):

Given a right triangle, one leg is 10 in. and the hypotenuse is 15 in., find the length of the other leg using the Pythagorean theorem. (Use sq rt for the square root symbol).

OpenStudy (pfenn1):

Do you know the Pythagorean Theorem?

OpenStudy (phoenixfire):

Pythagorean Theorem: \[y^2 = a^2 + b^2\] you know y and a. \[15^2=10^2+b^2\] rearrange for b: \[\sqrt{15^2-10^2} = b\]

OpenStudy (anonymous):

thx

OpenStudy (pfenn1):

you're welcome